Consider a solution containing 0.100 M fluoride ions and 0.126 M hydrogen fluoride. The concentration of fluoride ions after the addition of 4.00 mL of 0.0100 M HCl to 25.0 mL of this solution is __________ M.
Answer #1

initially

millimoles of fluoride ion = 25.0 x 0.100 = 2.50

millimoles of HCl added = 4.00 x 0.0100 = 0.04

after HCl added

millimoles of fluride ion = 2.50 - 0.04 = 2.46

total volume = 25.0 + 4.00 = 29.00 mL

[Fluoride] = 2.46 / 29.00

[fluoride] = 0.085 M